OpenStudy (anonymous):

PLZ HELP, what is the sum of 6/y+5 + 3/y-2 ?

OpenStudy (anonymous):

Can u help? @sami-21

OpenStudy (anonymous):

yes take LCM (y+5)(y-2)

OpenStudy (anonymous):

And do what

OpenStudy (anonymous):

this will become. \[\Large \frac{6(y-2)+5(y+5)}{(y+5)(y-2)}\]

OpenStudy (anonymous):

@sami-21 i don't get this! What's the answer,?

OpenStudy (anonymous):

denominator remains the same simplify numerator and tell me what you got .

OpenStudy (gw2011):

6/(y+5) + 3/)y-2) [6(y-2) + 3(y+5)]/(y+5)(y-2) (6y-12+3y+15)/(y+5)(y-2) (9y + 3)/(y+5)(y-2) 3(y+1)/(y+5)(y-2)

OpenStudy (anonymous):

So is the answer 9y+3/ (y+5)(y-2) ? @gw2011

OpenStudy (gw2011):

Yes, but 9y+3 can be simplified to 3(3y+1). Sorry, but I mistyped 9y+3=3(y+1). It should be 3(3y+1).

OpenStudy (anonymous):

Thankyou

OpenStudy (gw2011):

You're welcome. I'm glad I could help.